OpenStudy (anonymous):

Given an angle of 210 degrees, find the angle measure in radians and revolutions.

OpenStudy (amistre64):

2pi radians is equal to 360 degrees; so that can be used as a conversion rate and 1 rotation is equal to 2pi radians; or, 1 rotation is equal to 360 degrees can be used as conversion ratios for that part

OpenStudy (anonymous):

I know that, can you answer it though, and show me the steps?

OpenStudy (amistre64):

when given a conversion factor: a/b then some value: k, can be converted as the product: k a/b

OpenStudy (amistre64):

since i feel that you are adept at multiplication, ill leave that to you to decide the specific solutions ....

OpenStudy (anonymous):

Are you serious, Are you going to help me or what?

OpenStudy (amistre64):

do i have to get out my flash cards and we go over the multiplication tables?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

lol :) its going to take a while to get all the way up to the 210s

OpenStudy (anonymous):

It's*

OpenStudy (anonymous):

Maybe I should give you English lessons in exchange for math lessons. lol

OpenStudy (amistre64):

\[210(d)\frac{2pi(rads)}{360(d)}=p~(rads)\] \[210(d)\frac{1(rot)}{360(d)}=q~(rots)\]

OpenStudy (anonymous):

Set 2Pi equal to 180 degrees. You want to eliminate the unit in the problem to convert to the requested unit. So if you wanted to eliminate the "degree" to get radian, what would you do?

OpenStudy (amistre64):

its typing lessons i need, and less arthritis :)

OpenStudy (anonymous):

Math people are very concise. You answer was given is amistre64's first response.

OpenStudy (anonymous):

How do you find radians if the revolutions is 13/15?

OpenStudy (amistre64):

1 revolution is equal to 2pi radians

OpenStudy (anonymous):

do you mean to say 13/15 th of a degree? NOT revolution?

OpenStudy (anonymous):

I mean revolutions there is no 13/15 degree

OpenStudy (amistre64):

\[\frac{13}{15}(rev)\frac{2pi(rads)}{1(rev)}\]

OpenStudy (amistre64):

1 revolution is equal to 1 rotation .... so thats just semantics :)

OpenStudy (anonymous):

Are they multiplied or what?

OpenStudy (amistre64):

of course, which is what i expressed at the start ... when given a conversion factor: a/b then some value: k, can be converted as the product: k a/b

OpenStudy (amistre64):

then there was some schtick about multiplication tables and flash cards .... this all ties in with the big picture :)